# Weixiang Zhao1, Yulin Hu1   Jiahe Guo1footnotemark:    Xingyu Sui1footnotemark:   , Tongtong Wu2, Yang Deng3 Yanyan Zhao1, Bing Qin1, Wanxiang Che1, Ting Liu1 1

### v1

@@ Abstract
Despite the growing global demand for large language models (LLMs) that serve users from diverse linguistic backgrounds, most cutting-edge LLMs remain predominantly English-centric. This creates a performance gap across languages, restricting access to advanced AI services for non-English speakers. Current methods to enhance multilingual capabilities largely rely on data-driven post-training techniques, such as multilingual instruction tuning or continual pre-training. However, these approaches encounter significant challenges, including the scarcity of high-quality multilingual datasets and the limited enhancement of multilingual capabilities. They often suffer from off-target issues and catastrophic forgetting of central language abilities. To this end, we propose Lens, a novel approach to enhance multilingual capabilities of LLMs by leveraging their internal language representation spaces. Specially, Lens operates by manipulating the hidden representations within the language-agnostic and language-specific subspaces from top layers of LLMs. Using the central language as a pivot, the target language is drawn closer to it within the language-agnostic subspace, allowing it to inherit well-established semantic representations. Meanwhile, in the language-specific subspace, the representations of the target and central languages are pushed apart, enabling the target language to express itself distinctly. Extensive experiments on one English-centric and two multilingual LLMs demonstrate that Lens effectively improves multilingual performance without sacrificing the model’s original central language capabilities, achieving superior results with much fewer computational resources compared to existing post-training approaches.
@@ 1 Introduction
In an increasingly interconnected world, large language models (LLMs) are expected to cater to a diverse range of users across various linguistic backgrounds (Ouyang et al., 2023; Zhao et al., 2024a; Zheng et al., 2024). However, despite this global trend, most state-of-the-art LLMs remain predominantly English-centric (Brown et al., 2020; Touvron et al., 2023a; b; Jiang et al., 2023; AI@Meta, 2024). These models exhibit significantly better performance in English than in other languages, leading to an imbalance in user experience and potentially excluding large segments of the global population from accessing advanced AI services (Wang et al., 2024; Zhu et al., 2024b).
@@
This disparity has directly spurred research efforts to enhance the multilingual capabilities of LLMs, aiming to provide more equitable access and performance across various linguistic communities. Current approaches are predominantly based on data-driven post-training paradigm, such as multilingual instruction tuning (Zhang et al., 2023; Zhu et al., 2023; Üstün et al., 2024) or continual pre-training (Cui et al., 2023; Kuulmets et al., 2024; Jaavid et al., 2024), which primarily seeks to inject or elicit multilingual knowledge with the supervision signals from external datasets.
@@
While this paradigm is widely embraced and demonstrates certain successes, it faces several significant challenges. (1) The efficacy of multilingual enhancement heavily depend on large-scale and high-quality multilingual datasets (Zhou et al., 2023; Liu et al., 2024b), which are both time-consuming and labor-intensive to obtain for each language. (2) It favors improving multilingual understanding over generation capabilities, which leaves the off-target issue inadequately addressed. As a result, the model often struggles to generate accurate responses in the intended language when prompted (Zhang et al., 2020; Lai et al., 2024; Sennrich et al., 2024). (3) The model’s performance in languages it previously handled well is risking at catastrophic forgetting (McCloskey & Cohen, 1989), such as English in the LLaMA family (Touvron et al., 2023a; b; AI@Meta, 2024).
@@
In this work, we seek to provide a new perspective on addressing the aforementioned challenges by exploring and manipulating the internal representation within the language-related latent spaces of LLMs (Zou et al., 2023; Park et al., 2024). Taking the enhancement of multilingual capabilities for English-centric LLMs as an example. This is based on the intuitive idea that the well-established English representations in existing English-centric LLMs can act as a pivot to improve the performance of other languages. More specifically, for the target language to be extended, its language-agnostic semantic representations should be pulled close to those of English, enabling it to quickly inherit general abilities in English without the need for supervision signals from external multilingual training data. Conversely, the language-specific linguistic representations of the target language should be pushed away from English to avoid off-target issues, ensuring accurate responses in the target language. Also, during this process, it is crucial to ensure that the English pivot representation remains unchanged to effectively prevent catastrophic forgetting.
@@
To achieve this, we propose Lens, a novel multiLingual Enhancement method based on the hidden represeNtations within language Space of LLMs. To be more specific, Lens comprises two stages: Language Subspace Probing (LSP) and Language Subspace Manipulation (LSM). During LSP, the multilingual hidden space within a single layer of the backbone are decoupled into two orthogonal components: a language-agnostic subspace and a language-specific subspace. These subspaces are efficiently derived using singular value decomposition. Then in LSM, we align the parallel multilingual input representations of the target language and the central language in the language-agnostic subspace. This allows the target language to directly inherit the well-established semantic representations of the central language. Simultaneously, the projection components of the target language within the language-specific space are pushed away from those of the central language, guiding the target language toward its distinct linguistic expression and ensuring the target language is properly expressed thereby mitigating the off-target issue. Finally, we align the central language’s current representations with its original ones to preserve its proficiency during multilingual enhancement.
We conduct extensive experiments under bilingual and multilingual enhancement setups. Results on one English-centric (LLaMA-3-8B-Instruct) and 2 multilingual LLMs (LLaMA-3.1-8B-Instruct and Phi-3.5-mini-Instruct), demonstrate that Lens succeed to improve target languages on both multilingual comprehension and generation tasks without sacrificing the strong capability of central language, showing the efficacy and scalability of our method. Deeper analysis highlights the significance of steering the target language towards its unique expressions within its own language-specific subspace to fully enhance both comprehension and generation capabilities. This is overlooked by most existing approaches, which primarily focus on aligning representations across different languages to boost multilingual performance. It is crucial to note that, building on recent findings that language-related parameters are primarily concentrated in the top layers of LLMs (Wendler et al., 2024), Lens achieves high resource efficiency compared to baselines, with much fewer computational costs by only updating the model’s higher layers using just a few hundred data points.
@@
The main contributions of this work are summarized as follows: (1) We provide a novel perspective for the multilingual enhancement of large language models with their internal language representation space leveraged. (2) We propose Lens, an efficient and effective multilingual enhancement method that operates within the language representation space of large language models. (3) Extensive experiments on one English-centric and two multilingual LLMs demonstrate the effectiveness, efficiency, scalability of our method to obtain truly multilingual enhanced chat-style backbones without sacrificing original central language performance.
2 Related Works
Multilingual Large Language Model
With the acceleration of globalization, multilingual large language models (MLLMs) are gaining significant attention for their ability to handle multiple languages comprehensively (Qin et al., 2024). Pretraining on multilingual data is a common approach to gain the multilingual capabilities (Conneau & Lample, 2019; Xue et al., 2020; Lin et al., 2022; Shliazhko et al., 2022; Wei et al., 2023; Xue et al., 2022; Le Scao et al., 2023; Blevins et al., 2024). However, due to the uneven distribution of data in pretraining corpora, current LLMs or MLLMs exhibit uneven language capabilities, with most state-of-the-art models heavily biased towards English (Jiang et al., 2023; AI@Meta, 2024; Abdin et al., 2024). Moreover, pretraining from scratch is computationally intensive. These limitations have directly sparked research into expanding or enhancing the language capabilities of current LLMs or MLLMs.
Multilingual Enhancement for LLMs
The former focuses on leveraging the LLMs’ own translation capabilities to translate low-resource language inputs into the central language, and then generating a response (Shi et al., 2023; Huang et al., 2023; Qin et al., 2023; Etxaniz et al., 2024; Zhang et al., 2024b). For example, Huang et al. (2023) introduce cross-lingual-thought prompting to minimize language disparities. However, Liu et al. (2024a) reveal the limitations of these methods, showing they are not optimal for real-world scenarios and highlighting the necessity of more comprehensive multilingual enhancement.
The latter aims to conduct further multilingual post-training to inject or elicit extensive language knowledge for specific languages, including ways of continual pre-training (Zhang et al., 2021; Cui et al., 2023; Chen et al., 2023b; Lin et al., 2024; Kuulmets et al., 2024; Jaavid et al., 2024) and instruction tuning (Muennighoff et al., 2023; Chen et al., 2023c; Indurthi et al., 2024; Ahuja et al., 2024; Lai & Nissim, 2024; Zhang et al., 2024c; Zhu et al., 2024a; Li et al., 2024; Zhao et al., 2024c). For example, Cui et al. (2023) attempt to inject Chinese knowledge into LLaMA by conducting continual pre-training on a large-scale Chinese corpus, while Zhu et al. (2023) focus more on building language alignment through cross-lingual instruction tuning and translation training.
Our proposed Lens stands out from existing methods in that we seek multilingual supervision signals from the internal language representation space of the LLMs, rather than relying primarily on external multilingual datasets as in the above methods, which offers fresh insights and new opportunities for enhancing the multilingual capabilities of LLMs both efficiently and effectively.
Representation Engineering
Editing or manipulating representation within LLMs has garnered increasing attention due to its transparency and lightweight properties (Zou et al., 2023). This is theoritically rooted from Linear Representation Hypothesis (Mikolov et al., 2013; Nanda et al., 2023; Park et al., 2024), which posits that various human-interpretable concepts are encoded in linear subspaces of model representations. Building upon this, exist works attempt to edit representations at inference time to develop models that are more truthful (Li et al., 2023b; Campbell et al., 2023; Zhang et al., 2024a), and harmless (Lee et al., 2024; Uppaal et al., 2024). We expand and implement this paradigm for the multilingual enhancement of LLMs by focusing on representations during the training phase, ensuring that the efficiency of LLMs remains unaffected during the inference phase.
@@ 3.2 Language Subspace Probing
$\bm{m}_{l}=\frac{1}{n}\sum_{i=1}^{n}\bm{e}_{l}^{i}$ (1)
where $\bm{e}_{l}^{i}\in\mathbb{R}^{d}$ is the embedding of the last token for the $i$-th sample in language $l$, and $n$ is the total number of samples for each language. Concatenating $\bm{m}_{l}$ of $L$ languages column-by-column results in the mean embedding matrix $\bm{M}\in\mathbb{R}^{d\times L}$ specifying the multilingual latent space.
Follow previous works (Pires et al., 2019; Libovickỳ et al., 2020; Yang et al., 2021), we hypothesize that such multilingual latent space $M$ could be decomposed into two orthogonal components (1) a language-agnostic subspace $\bm{M}_{a}$ representing what is commonly shared across languages and (2) a language-specific one $\bm{M}_{s}$ specifying on which different languages express different linguistic signals. Following Piratla et al. (2020); Xie et al. (2022), the objective can be formulated as:
$\min_{\bm{M}_{a},\bm{M}_{s},\bm{\Gamma}}$ $\left\|\bm{M}-\bm{M}_{a}\bm{\mathbbm{1}}^{\top}-\bm{M}_{s}\bm{\Gamma}^{\top}\right\|_{F}^{2}$ (2)
s.t. $\text{Span}\left(\bm{M}_{a}\right)\perp\text{Span}\left(\bm{M}_{s}\right),$
where $\bm{M}_{a}\in\mathbb{R}^{d\times 1}$, $\bm{M}_{s}\in\mathbb{R}^{d\times r}$ and $\bm{\Gamma}\in\mathbb{R}^{L\times r}$ is the coordinates of language-specific signals along the subspace’s $r$ components. And a lower dimensionality for $\bm{M}_{a}$ is reasonable because the semantic consistency across different languages can be captured in a simpler form. Meanwhile, $\bm{M}_{s}$ requires a higher dimensionality to account for the distinct features of each language.
The optimal solution of Equation 2 can be computed efficiently via Singular Value Decomposition (SVD), where Algorithm 1 in Appendix B presents the detailed procedure.
After obtaining the language-specific subspace $\bm{M}_{s}$, we aim to identify a direction of language expression within this subspace, which points from the projection of mean representation from target language $\bm{m}_{l}$ to that from central language $\bm{m}_{c}$. Formally, the linguistic language expression direction $\bm{\delta}_{l}\in\mathbb{R}^{d}$ for each target language $l$ is calculated as:
$\bm{\delta}_{l}=\bm{M}_{s}^{T}\bm{M}_{s}(\bm{m}_{l}-\bm{m}_{c})$ (3)
Figure 1: The overall architecture of our proposed Lens for multilingual enhancement. (1) In the LSP, we begin by decomposing the multilingual latent space, which is formed by the representations of probing samples from both the target and central languages. Using singular value decomposition (SVD), we separate this space into two orthogonal components: a language-agnostic subspace, $\bm{M}_{a}$, and a language-specific subspace, $\bm{M}_{s}$. (2) Then in LSM, the parallel multilingual representations of the target languages are pushed toward their respective linguistic expression directions within $\bm{M}_{s}$, while being pulled closer to the central language in $\bm{M}_{a}$. Additionally, the representations of the central language are carefully constrained to remain largely intact.
@@ 3.3 Language Subspace Manipulation
First, We propose to pull parallel multilingual representations closer within the shared language-agnostic subspace $\bm{M}_{a}$. This allows us to directly inherit the well-established general capabilities of the central language. Formally, this goal is accomplished by projecting multilingual representations (at the position of the last token) onto the subspace $\bm{M}_{a}$, with the optimization objective defined as:
$\mathcal{L}_{1}=\left\|\bm{M}_{a}^{T}\bm{M}_{a}(\bm{x}_{l}-\bm{x}_{c})\right\|^{2}$ (4)
where $\bm{x}_{l}$ and $\bm{x}_{c}$ are parallel multilingual representations from target language $l$ and central one.
Second, to ensure that each target language can be accurately expressed and to alleviate the off-target issue, we need to push the multilingual representations in the language-specific subspace $\bm{M}_{s}$ towards their respective language-specific expression directions. This can be achieved through the projection onto the subspace $\bm{M}_{s}$ and optimizing the following objective:
$\mathcal{L}_{2}=\left\|\bm{M}_{s}^{T}\bm{M}_{s}(\bm{x}_{l}-\bm{x}_{l}^{\text{ref}})-\lambda_{l}\bm{\delta}_{l}\right\|^{2}$ (5)
where $\bm{x}_{l}^{\text{ref}}$ is the representation of target language $l$ obtained from original reference model and $\lambda_{l}$ is a scalar of push strength for the corresponding language. The above process can be interpreted as directing the language-specific representations of each target language to shift a specific distance from their original positions toward a direction that enables accurate expression.
@@
$\mathcal{L}_{3}=\left\|\bm{x}_{c}-\bm{x}_{c}^{\text{ref}}\right\|^{2}$ (6)
where $\bm{x}_{c}^{\text{ref}}$ is the representation of central language $c$ obtained from original reference model.

### v2

# for Large Language Models

@@ Abstract
As global demand for multilingual large language models (LLMs) grows, most LLMs still remain overly focused on English, leading to the limited access to advanced AI for non-English speakers. Current methods to enhance multilingual capabilities largely rely on data-driven post-training techniques, such as multilingual instruction tuning or continual pre-training. However, these approaches exhibit significant limitations, including high resource cost, exacerbation of off-target issue and catastrophic forgetting of central language abilities. To this end, we propose Lens, a novel approach that enhances multilingual capabilities by leveraging LLMs’ internal language representation spaces. Lens operates on two subspaces: the language-agnostic subspace, where it aligns target languages with the central language to inherit strong semantic representations, and the language-specific subspace, where it separates target and central languages to preserve linguistic specificity. Experiments on three English-centric LLMs show that Lens significantly improves multilingual performance while maintaining the model’s English proficiency, achieving better results with less computational cost compared to existing post-training approaches. Our code is available at https://github.com/circle-hit/Lens.
Lens: Rethinking Multilingual Enhancement
for Large Language Models
Weixiang Zhao1, Yulin Hu1thanks:   Equal contribution  Jiahe Guo1footnotemark:  Xingyu Sui1footnotemark: , Tongtong Wu2 Yang Deng3 Yanyan Zhao1, Bing Qin1, Wanxiang Che1, Ting Liu1 1Harbin Institute of Technology, 2Monash University 3Singapore Management University {wxzhao,yyzhao,qinb,car,tliu}@ir.hit.edu.cn
@@ 1 Introduction
In an increasingly interconnected world, large language models (LLMs) are expected to cater to a diverse range of users across various linguistic backgrounds (Ouyang et al., 2023; Zhao et al., 2024a; Zheng et al., 2024; Qin et al., 2024). However, most state-of-the-art LLMs remain heavily English-centric (Brown et al., 2020; Touvron et al., 2023a, b; Jiang et al., 2023; AI@Meta, 2024), performing far better in English than other languages, thereby potentially marginalizing large portions of the global population from accessing advanced AI services (Wang et al., 2024a; Zhu et al., 2024b).
@@
This disparity has directly spurred research efforts to enhance multilingual capabilities of LLMs. Current approaches are predominantly based on data-driven post-training paradigm, such as multilingual instruction tuning (Zhang et al., 2023; Zhu et al., 2023; Üstün et al., 2024) or continual pre-training (Cui et al., 2023; Kuulmets et al., 2024; Jaavid et al., 2024), which primarily seeks to elicit cross-lingual alignment (Schuster et al., 2019) or inject multilingual knowledge with the supervision signals from external datasets.
@@
While this paradigm is widely embraced and demonstrates certain successes, it faces several significant limitations. (1) The efficacy often depends on training with large-scale multilingual datasets (Cui et al., 2023; Zhu et al., 2023), which incur large computational overhead. (2) It overly emphasizes alignment across languages, neglecting the modeling of language-specific features, which exacerbates the off-target issue (Zhang et al., 2020). As a result, the model often struggles to generate accurate responses in the intended language when prompted (Lai et al., 2024; Sennrich et al., 2024). (3) The model’s performance in languages it previously handled well is risking at catastrophic forgetting (McCloskey and Cohen, 1989).
@@
In this work, we seek to provide a new perspective on addressing the aforementioned limitations by exploring and manipulating the internal representation within the language-related latent spaces of LLMs (Zou et al., 2023; Park et al., 2024). Taking the enhancement of multilingual capabilities for English-centric LLMs as an example. This is based on the intuitive idea that the well-established English representations in existing English-centric LLMs can act as a pivot to improve the performance of other languages. More specifically, for the target language to be enhanced, its language-agnostic semantic representations should be pulled close to those of English, facilitating cross-lingual alignment. Meanwhile, the language-specific linguistic representations should be pushed away from English to preserve the unique representation of each language. Also, during this process, it is crucial to ensure that the English pivot representation remains unchanged to effectively prevent catastrophic forgetting.
Figure 1: The overall architecture of our proposed Lens for multilingual enhancement. (1) In the LSP, we begin by decomposing the multilingual latent space, which is formed by the representations of probing samples from both the target and central languages. Using singular value decomposition (SVD), we separate this space into two orthogonal components: a language-agnostic subspace, $\boldsymbol{M}_{a}$, and a language-specific subspace, $\boldsymbol{M}_{s}$. (2) Then in LSM, the parallel multilingual representations of the target languages are pushed toward their respective linguistic expression directions within $\boldsymbol{M}_{s}$, while being pulled closer to the central language in $\boldsymbol{M}_{a}$. Additionally, the representations of the central language are carefully constrained to remain largely intact.
@@
To achieve this, we propose Lens, a novel multiLingual Enhancement method based on the hidden represeNtations within language Space of LLMs. To be more specific, Lens comprises two stages: Language Subspace Probing (LSP) and Language Subspace Manipulation (LSM). During LSP, the multilingual hidden space within a single layer of the backbone are decoupled into two orthogonal components, a language-agnostic subspace and a language-specific subspace, via singular value decomposition. Then in LSM, we align the parallel multilingual input representations of the target language and the central language in the language-agnostic subspace. Simultaneously, the projection components of the target language within the language-specific space are pushed away from those of the central language, guiding the target language toward its distinct linguistic expression and ensuring the target language is properly expressed thereby mitigating the off-target issue. Finally, we align the central language’s current representations with its original ones to preserve its proficiency during multilingual enhancement. It is crucial to note that, building on recent findings that language-related parameters are primarily concentrated in the top layers of LLMs (Wendler et al., 2024), Lens only updates the higher layers of the backbone with just a few hundred data points, exhibiting high resource efficiency.
We conduct extensive experiments on bilingual and multilingual enhancement setups, targeting languages from diverse linguistic families and varying resource levels. Results on three English-centric LLMs (LLaMA-3-8B-Instruct, LLaMA-3.1-8B-Instruct and Phi-3.5-mini-Instruct) show that Lens outperform baselines and open-source multilingual-enhanced LLMs on both multilingual comprehension and generation tasks. Notably, Lens achieves these improvements without compromising the central language’s strong capabilities, while requiring significantly less computational overhead, highlighting its effectiveness, efficiency, and scalability.
@@
The main contributions of this work are summarized as follows: (1) We provide a novel perspective for the multilingual enhancement of LLMs with their internal language representation space leveraged. (2) We propose Lens, an efficient and effective multilingual enhancement method that operates within the language representation space of large language models. (3) Extensive experiments demonstrate the effectiveness, efficiency, scalability of our method to obtain multilingual enhanced chat-style backbones without sacrificing original central language performance.
@@ 2.2 Language Subspace Probing
$\boldsymbol{m}_{l}=\frac{1}{n}\sum_{i=1}^{n}\boldsymbol{e}_{l}^{i}$ (1)
where $\boldsymbol{e}_{l}^{i}\in\mathbb{R}^{d}$ is the embedding of the last token for the $i$-th sample in language $l$, and $n$ is the total number of samples for each language. Concatenating $\boldsymbol{m}_{l}$ of $L$ languages column-by-column results in the mean embedding matrix $\boldsymbol{M}\in\mathbb{R}^{d\times L}$ specifying the multilingual latent space.
Following previous works (Pires et al., 2019; Libovickỳ et al., 2020; Yang et al., 2021), we hypothesize that such multilingual latent space $M$ could be decomposed into two orthogonal components (1) a language-agnostic subspace $\boldsymbol{M}_{a}$ representing what is commonly shared across languages and (2) a language-specific one $\boldsymbol{M}_{s}$ specifying on which different languages express different linguistic signals. Following Piratla et al. (2020); Xie et al. (2022), the objective can be formulated as:
$\min_{\boldsymbol{M}_{a},\boldsymbol{M}_{s},\boldsymbol{\Gamma}}$ $\left\|\boldsymbol{M}-\boldsymbol{M}_{a}\boldsymbol{\mathbbm{1}}^{\top}-\boldsymbol{M}_{s}\boldsymbol{\Gamma}^{\top}\right\|_{F}^{2}$ (2)
s.t. $\text{Span}\left(\boldsymbol{M}_{a}\right)\perp\text{Span}\left(\boldsymbol{M}_{s}\right),$
where $\boldsymbol{M}_{a}\in\mathbb{R}^{d\times 1}$, $\boldsymbol{M}_{s}\in\mathbb{R}^{d\times r}$ and $\boldsymbol{\Gamma}\in\mathbb{R}^{L\times r}$ is the coordinates of language-specific signals along the subspace’s $r$ components. And a lower dimensionality for $\boldsymbol{M}_{a}$ is reasonable because the semantic consistency across different languages can be captured in a simpler form. Meanwhile, $\boldsymbol{M}_{s}$ requires a higher dimensionality to account for the distinct features of each language.
The optimal solution of Equation 2 can be computed efficiently via Singular Value Decomposition (SVD), where Algorithm 1 in Appendix A presents the detailed procedure.
After obtaining the language-specific subspace $\boldsymbol{M}_{s}$, we aim to identify a direction of language expression within this subspace, which points from the projection of mean representation from target language $\boldsymbol{m}_{l}$ to that from central language $\boldsymbol{m}_{c}$. Formally, the linguistic language expression direction $\boldsymbol{\delta}_{l}\in\mathbb{R}^{d}$ for each target language $l$ is calculated as:
$\boldsymbol{\delta}_{l}=\boldsymbol{M}_{s}^{T}\boldsymbol{M}_{s}(\boldsymbol{m}_{l}-\boldsymbol{m}_{c})$ (3)
@@ 2.3 Language Subspace Manipulation
First, we propose to pull parallel multilingual representations closer within the shared language-agnostic subspace $\boldsymbol{M}_{a}$. This allows us to directly inherit the well-established general capabilities of the central language. Formally, this goal is accomplished by projecting multilingual representations (at the position of the last token) onto the subspace $\boldsymbol{M}_{a}$, with the optimization objective defined as:
$\mathcal{L}_{1}=\left\|\boldsymbol{M}_{a}^{T}\boldsymbol{M}_{a}(\boldsymbol{x}_{l}-\boldsymbol{x}_{c})\right\|^{2}$ (4)
where $\boldsymbol{x}_{l}$ and $\boldsymbol{x}_{c}$ are parallel multilingual representations from target language $l$ and central one.
Second, to ensure that each target language can be accurately expressed and to alleviate the off-target issue, we need to push the multilingual representations in the language-specific subspace $\boldsymbol{M}_{s}$ towards their respective language-specific expression directions. This can be achieved through the projection onto the subspace $\boldsymbol{M}_{s}$ and optimizing the following objective:
$\mathcal{L}_{2}=\left\|\boldsymbol{M}_{s}^{T}\boldsymbol{M}_{s}(\boldsymbol{x}_{l}-\boldsymbol{x}_{l}^{\text{ref}})-\lambda_{l}\boldsymbol{\delta}_{l}\right\|^{2}$ (5)
where $\boldsymbol{x}_{l}^{\text{ref}}$ is the representation of target language $l$ obtained from original reference model and $\lambda_{l}$ is a scalar of push strength for the corresponding language. The above process can be interpreted as directing the language-specific representations of each target language to shift a specific distance from their original positions toward a direction that enables accurate expression.
@@
$\mathcal{L}_{3}=\left\|\boldsymbol{x}_{c}-\boldsymbol{x}_{c}^{\text{ref}}\right\|^{2}$ (6)
where $\boldsymbol{x}_{c}^{\text{ref}}$ is the representation of central language $c$ obtained from original reference model.